Product Decisions Under Bottlenecked Operations Youngstown Glass Company manufactures three types of safety plate glass: large,...

  1. Product Decisions Under Bottlenecked Operations

    Youngstown Glass Company manufactures three types of safety plate glass: large, medium, and small. All three products have high demand. Thus, Youngstown Glass is able to sell all the safety glass that it can make. The production process includes an autoclave operation, which is a pressurized heat treatment. The autoclave is a A condition that occurs when product demand exceeds production capacity.production bottleneck. Total fixed costs are $234,000 for the company as a whole. In addition, the following information is available about the three products:

    Large Medium Small
    Unit selling price $122 $314 $158
    Unit variable cost (96) (257) (139)
    Unit contribution margin $ 26 $ 57 $ 19
    Autoclave hours per unit 4 6 2
    Total process hours per unit 12 18 4
    Budgeted units of production 5,100 5,100 5,100

    a. Determine the contribution margin by glass type and the total company operating income for the budgeted units of production.

    Large      Medium      Small      Total     
    Units produced
    Revenues $ $ $ $
    Variable costs
    Contribution margin $ $ $ $
    Fixed costs
    Operating income $

    b. Prepare an analysis showing which product is the most profitable per bottleneck hour. Round the "Unit contribution margin per production bottleneck hour" amounts to the nearest cent.

    Large     Medium     Small    
    Contribution margin $ $ $
    Autoclave hours per unit
    Unit contribution margin per production bottleneck hour $ $ $
Answer #1
a
Large    Medium Small Total
Units produced 5100 5100 5100
Revenues 622200 1601400 805800 3029400
Variable costs 489600 1310700 708900 2509200
Contribution margin 132600 290700 96900 520200
Fixed costs 234000
Operating income 286200
b
Large    Medium Small
Contribution margin 26 57 19
Autoclave hours per unit 4 6 2
Unit contribution margin per production bottleneck hour 6.50 9.50 9.50
Workings:
Large    Medium Small
Revenues =5100*122 =5100*314 =5100*158
Variable costs =5100*96 =5100*257 =5100*139